# Fading ergodicity

### v1

@@ Abstract
Eigenstate thermalization hypothesis (ETH) represents a breakthrough in many-body physics since it allows to link thermalization of physical observables with the applicability of random matrix theory (RMT). Recent years were also extremely fruitful in exploring possible counterexamples to thermalization, ranging, among others, from integrability, single-particle chaos, many-body localization, many-body scars, to Hilbert-space fragmentation. In all these cases the conventional ETH is violated. However, it remains elusive how the conventional ETH breaks down when one approaches the boundaries of ergodicity, and whether the range of validity of the conventional ETH coincides with the validity of RMT-like spectral statistics. Here we bridge this gap and we introduce a scenario of the ETH breakdown in many-body quantum systems, which establishes a link between the conventional ETH and non-ergodic behavior. We conjecture this scenario to be relevant for the description of finite many-body systems at the boundaries of ergodicity, and we provide numerical and analytical arguments for its validity in the quantum sun model of ergodicity breaking phase transition. For the latter, we provide evidence that the breakdown of the conventional ETH is not associated with the breakdown of the RMT-like spectral statistics.
Introduction. The quantum chaos conjecture links the emergence of random-matrix theory (RMT) statistics in quantum systems with the chaotic dynamics in their classical limit bohigas_giannoni_84; casati_valzgris_80. Perhaps surprisingly or not, the RMT predictions are also relevant for the description of spectral statistics of quantum many-body systems without classical counterparts, e.g., of interacting spin-1/2 systems on a lattice montambaux_poilblanc_93; hsu_dauriac_93; distasio_zotos_95; prosen_99; santos_04; rabson_narozhny_04; rigol_santos_10. Nevertheless, the experiments that study nonequilibrium dynamics of isolated quantum many-body systems usually cannot access the spectral properties, but they can measure local observables such as site occupations bloch_dalibard_12. Here, the central role is played by the eigenstate thermalization hypothesis (ETH) deutsch_91; srednicki_94; rigol_dunjko_08, which provides simple principles to explain the agreement between the observable expectation values in time-evolving pure states and the predictions of statistical ensembles dalessio_kafri_16.
@@
When considering counterexamples to thermalization, the ETH from Eq. (1) is not expected to be valid. Specifically, two features of the matrix elements may emerge in non-ergodic systems: (a) the fluctuations of matrix elements may decay polynomially (instead of exponentially) with $L$, and (b) the matrix elements in some eigenstates, dubbed outliers, may not approach the corresponding microcanonical averages. These frameworks allow for introducing weaker forms of ETH that apply, e.g., for integrable systems biroli_lauchli_10; ikeda_ueda_15; alba_15; leblond_mallayya_19; brenes_leblond_20, single-particle chaotic systems lydzba_swietek_24, many-body localization luitz_16; panda_scardicchio_19; Colmenarez2019; corps_molina_21, many-body scars shiraishi_mori_17; *mondaini_mallayya_18; serbyn_abanin_2021; moudgalya_2022, and Hilbert-space fragmentation sala_rakovszky_2020; moudgalya_2022. However, all these forms of ETH are incompatible with ergodicity and thermalization. These considerations give rise to the central question of our study: how and when does the conventional ETH from Eq. (1) evolve into other (weaker) forms of ETH when ergodicity is fading, and to what extent is this transition related to the breakdown of RMT-like short-range spectral statistics?
@@
Here, our goal is to go beyond the paradigm of ergodicity breaking as the absence of validity of RMT spectral statistics, and hence we focus our analysis on observables. We introduce a scenario in which the fluctuations of the observable matrix elements soften when approaching the ergodicity boundary, even though the observables still thermalize and the short-range spectral statistics comply with RMT predictions. Our theory provides a natural bridge between the two limits, i.e., the conventional ETH limit (1) and the completely non-ergodic limit. Moreover, in the regime intermediate to these two limits, it establishes ergodicity beyond the conventional ETH. Hence, it can be used as observable-based precursor of the ergodicity breaking phase transition.
Scenarios for the breakdown of ETH. We now discuss the possible scenarios for the ETH breakdown in many-body quantum systems. The latter should be searched for observables that are diagonal in the basis in which the Hamiltonian eigenstates in the non-ergodic regime exhibit signatures of localization. This perspective is supported by the recent analysis of the matrix elements of observables in single-particle eigenstates of an Anderson insulator lydzba_zhang_21. As the limiting non-ergodic behavior, we hence expect the majority of the weight of the matrix elements to be accumulated in the diagonal matrix elements. An important technical aspect when considering ETH is the existence of a sum rule for the matrix elements of observables lydzba_swietek_24, which we impose as
@@
where $\cal D$ is the Hilbert space dimension and $O_{nm}\equiv\langle n|\hat{O}|m\rangle$, see SM for details.
@@
We note that recent studies based on the norms of adiabatic gauge potentials contributed valuable insights into the structure of the off-diagonal matrix elements in the vicinity of the ergodicity breakdown pandey_claeys_20. In particular, they suggested that the softening of fluctuations of matrix elements at low $\omega$ is s smooth process in the vicinity of the breakdown leblond_sels_21. However, a clear distinction between the scenarios (a) and (b) has to our knowledge not yet been established.
Here we put forth arguments in favor of scenario (b). Our main conjecture is that the fluctuations of the diagonal and low-$\omega$ off-diagonal matrix elements soften when approaching the ergodicity boundary, such that the fluctuating part of the ETH ansatz in Eq. (1) should acquire $\omega$-dependence, $\rho(\bar{E})^{-1/2}\to\Sigma(\bar{E},\omega)$. Hence, the difference of our scenario with respect to the conventional ETH is that the $L$-dependence of fluctuations, expressed via the density of states $\rho(\bar{E})$, and the $\omega$-dependence of the off-diagonal matrix elements, do not decouple as in Eq. (1). In the low-$\omega$ regime $\omega\to 0$, we express the softening of fluctuations as
@@
with $\eta\geq 2$. We propose that $\eta$ is determined by the ratio of the Thouless and Heisenberg energy, and when the Thouless time is sufficiently large, the divergence of $\eta$ signals proximity of the ergodicity breaking phase transition. This scenario is sketched in Fig. 1.
@@
Softening of ETH at small energies. We now present an argument supporting Eq. (3), i.e., a scenario that naturally interpolates between the conventional ETH ($\eta=2$) and its breakdown in the non-ergodic limit ($\eta\to\infty$). The argument is based on the $L$-dependence of the off-diagonal matrix elements at low $\omega$ for the targeted observable $\hat{O}$. For simplicity, we model the coarse-grained off-diagonal matrix elements at low $\omega$ with a Lorentzian function. This specific functional form is not crucial for the derivation of our main result, however, since we are approaching the ergodicity boundary from the ergodic side, it is important that the function is a constant in the $\omega\to 0$ limit. We consider
@@
Importantly, the Thouless energy $\Gamma$ depends on $L$ in most physical systems. Consequently, also the scaled off-diagonal matrix elements $\overline{|O_{nm}|^{2}}\rho$ are expected to depend on $L$. This is a known property, which is usually interpreted via the $L$-dependence of $f(\bar{E},\omega)$, see Eq. (1), at $\omega\ll\Gamma$ dalessio_kafri_16; luitz_barlev_16; brenes_leblond_20; richter_dymarsky_20; schoenle_jansen_21; wang_lamann_22; Sugimoto2022; balachandran_santos_23. However, in many cases this $L$-dependence is polynomial, and it is hence subleading when compared to the exponentially increasing $\rho$.
@@
Example: Quantum sun model. We now present analytical and numerical evidence that supports the above scenario in a toy model that hosts an ergodicity breaking phase transition in the thermodynamic limit, i.e., the quantum sun model of the avalanche theory deroeck_huveneers_17; luitz_huveneers_17; suntajs_vidmar_22. The Hamiltonian is defined as
$\hat{H}=\hat{H}_{\rm dot}+\sum_{j=1}^{L}\alpha^{u_{j}}\hat{S}_{n_{j}}^{x}\hat{S}_{j}^{x}+\sum_{j=1}^{L}h_{j}\hat{S}_{j}^{z},$ (7)
where $\hat{H}_{\rm dot}$ is a $2^{N}\times 2^{N}$ random matrix drawn from the Gaussian orthogonal ensemble (GOE) describing all-to-all interacting particles within an ergodic quantum dot (we set $N=3$). The second term defines the coupling between a spin $j$ outside the dot ($j=1,...,L)$ and a randomly selected spin $n_{j}$ within the dot, with $\alpha$ acting as the tuning parameter of the ergodicity breaking phase transition, and $u_{j}\propto j$. Details of the model implementation and the choice of model parameters are given in SM.
@@
In Fig. 2 and 3 we numerically test the above predictions. We focus on the observable $\hat{O}=\hat{S}_{L}^{z}$, i.e., the projection of the most weakly coupled spin on the z-axis. The physical motivation for selecting this observable is based on understanding that observables, which measure properties of spins that are most weakly coupled to the ergodic quantum dot, are most sensitive to ergodicity breaking suntajs_hopjan_24. Results for other observables are shown in SM.
@@
In Fig. 2, we show the $\omega$-dependence of the coarse-grained off-diagonal matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}$, where $(S_{L}^{z})_{nm}\equiv\langle n|\hat{S}_{L}^{z}|m\rangle$ and the coarse graining is carried out over a narrow window above the target $\omega$. The scaled matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}\rho$ vs $\omega$ exhibit accumulation of spectral weight at low $\omega$ and its depletion at large $\omega$, see the arrows in the main panel of Fig. 2. Remarkably, the results in Fig. 2 are shown for $\alpha=0.86$, for which the nearest level-spacing statistics complies with the GOE predictions SM. This suggests that the softening of the fluctuations of the matrix elements emerges while the short-range statistics are still GOE-like. Another perspective on the softening of the low-$\omega$ matrix elements is shown in the inset of Fig. 2, in which we show the scaled matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}\rho\Gamma$ vs $\omega/\Gamma$, exhibiting a reasonably good collapse of the results for all system sizes under investigation.
@@
Figure 3: Scaling of fluctuations of matrix elements. (a), (b) Eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, $\delta_{n}\equiv|(\hat{S}^{z}_{L})_{n+1}-(\hat{S}^{z}_{L})_{n}|$. We show the average and the maximal outliers in (a) and (b), respectively, see also SM. (c) Low-$\omega$ off-diagonal matrix elements, taken from a narrow energy window around the target $\omega=\sqrt{\Gamma\Delta}$. Dashed lines in (a)-(c) are the two-parameter fits of the function $a_{0}2^{-L/\eta}$ with the values of $\eta$ shown next to them. (d) Fluctuation exponents $\eta$ from Eq. (3) as a function of $\alpha$. The solid line is a fit of $b_{0}\eta^{*}$ to the results for the off-diagonal matrix elements, where $b_{0}=1.12$ is the fitting parameter and $\eta^{*}$ is the function from Eq. (10), in which $\tilde{\alpha}_{c}$ is replaced by a fitting parameter $\alpha_{c}$. We get $\alpha_{c}=0.745$, see the vertical dashed-dotted line, which is very close to the values of transition point obtained from other indicators such as the average gap ratio SM.
@@
A quantitative analysis of the fluctuations of the matrix elements in carried out in Fig. 3. We study the average eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, see Fig. 3(a), and the fluctuations of the low-$\omega$ off-diagonal matrix elements measured at $\omega=\sqrt{\Delta\Gamma}$, see Fig. 3(c). They both exhibit a fast decay deep in the ergodic regime (at $\alpha$ close to $\alpha=1$), while the decay becomes much slower when the ergodicity breaking transition point is approached, $\alpha\to\alpha_{c}$. The extracted fluctuation exponents $\eta$ as functions of $\alpha$ are shown as symbols in Fig. 3(d). They behave very similarly for the diagonal and the low-$\omega$ off-diagonal matrix elements, and remarkably, are well described by the solid line that corresponds to the analytical prediction from Eq. (10). Moreover, Fig. 3(b) shows that even the maximal outliers of the diagonal matrix elements decay to zero with increasing $L$. This property suggest that the system is still ergodic, and makes a clear distinction to non-ergodic systems such as integrable interacting systems in which the maximal outliers do not decay with system size biroli_lauchli_10; leblond_mallayya_19; lydzba_mierzejewski_23. In SM we also show the distributions of the matrix elements. They appear to be close to a normal distribution in nearly entire ergodic phase.

### v2

@@ Abstract
Eigenstate thermalization hypothesis (ETH) represents a breakthrough in many-body physics since it allows to link thermalization of physical observables with the applicability of random matrix theory (RMT). Recent years were also extremely fruitful in exploring possible counterexamples to thermalization, ranging, among others, from integrability, single-particle chaos, many-body localization, many-body scars, to Hilbert-space fragmentation. In all these cases the conventional ETH is violated. However, it remains elusive how the conventional ETH breaks down when one approaches the boundaries of ergodicity, and whether the range of validity of the conventional ETH coincides with the validity of RMT-like spectral statistics. Here we bridge this gap and we introduce a scenario of the ETH breakdown in many-body quantum systems, dubbed fading ergodicity regime, which establishes a link between the conventional ETH and non-ergodic behavior. We conjecture this scenario to be relevant for the description of finite many-body systems at the boundaries of ergodicity, and we provide numerical and analytical arguments for its validity in the quantum sun model of ergodicity breaking phase transition. For the latter, we provide evidence that the breakdown of the conventional ETH is not associated with the breakdown of the RMT-like spectral statistics.
I Introduction
The quantum chaos conjecture links the emergence of random-matrix theory (RMT) statistics in quantum systems with the chaotic dynamics in their classical limit bohigas_giannoni_84 ; casati_valzgris_80 . Perhaps surprisingly or not, the RMT predictions are also relevant for the description of spectral statistics of quantum many-body systems without classical counterparts, e.g., of interacting spin-1/2 systems on a lattice montambaux_poilblanc_93 ; hsu_dauriac_93 ; distasio_zotos_95 ; prosen_99 ; santos_04 ; rabson_narozhny_04 ; rigol_santos_10 . Nevertheless, the experiments that study nonequilibrium dynamics of isolated quantum many-body systems usually cannot access the spectral properties, but they can measure local observables such as site occupations bloch_dalibard_12 . Here, the central role is played by the eigenstate thermalization hypothesis (ETH) deutsch_91 ; srednicki_94 ; rigol_dunjko_08 , which provides simple principles to explain the agreement between the observable expectation values in time-evolving pure states and the predictions of statistical ensembles dalessio_kafri_16 .
@@ I Introduction
When considering counterexamples to thermalization, the ETH from Eq. (1) is not expected to be valid. Specifically, two features of the matrix elements may emerge in non-ergodic systems: (a) the fluctuations of matrix elements may decay polynomially (instead of exponentially) with $L$, and (b) the matrix elements in some eigenstates, dubbed outliers, may not approach the corresponding microcanonical averages. These frameworks allow for introducing weaker forms of ETH that apply, e.g., for integrable systems biroli_lauchli_10 ; ikeda_ueda_15 ; alba_15 ; leblond_mallayya_19 ; brenes_leblond_20 , single-particle chaotic systems lydzba_swietek_24 , many-body localization luitz_16 ; panda_scardicchio_19 ; Colmenarez2019 ; corps_molina_21 , many-body scars shiraishi_mori_17 ; *mondaini_mallayya_18; serbyn_abanin_2021 ; moudgalya_2022 , and Hilbert-space fragmentation sala_rakovszky_2020 ; khemani_hermele_2020 ; moudgalya_2022 . However, all these forms of ETH are incompatible with ergodicity and thermalization. These considerations give rise to the central question of our study: how and when does the conventional ETH from Eq. (1) evolve into other (weaker) forms of ETH when ergodicity is fading, and to what extent is this transition related to the breakdown of RMT-like short-range spectral statistics?
@@
Here, our goal is to go beyond the paradigm of ergodicity breaking as the absence of validity of RMT spectral statistics, and hence we focus our analysis on observables. We introduce a scenario in which the fluctuations of the observable matrix elements soften when approaching the ergodicity boundary, even though the observables still thermalize and the short-range spectral statistics comply with RMT predictions. Our theory provides a natural bridge between the two limits, i.e., the conventional ETH limit (1) and the completely non-ergodic limit. Moreover, in the regime intermediate to these two limits, dubbed fading ergodicity regime, it establishes ergodicity beyond the conventional ETH. Hence, it can be used as observable-based precursor of the ergodicity breaking phase transition.
II Scenarios for the ETH breakdown
We now discuss the possible scenarios for the ETH breakdown in many-body quantum systems. The latter should be searched for observables that are diagonal in the basis in which the Hamiltonian eigenstates in the non-ergodic regime exhibit signatures of localization. This perspective is supported by the recent analysis of the matrix elements of observables in single-particle eigenstates of an Anderson insulator lydzba_zhang_21 . As the limiting non-ergodic behavior, we hence expect the majority of the weight of the matrix elements to be accumulated in the diagonal matrix elements. An important technical aspect when considering ETH is the existence of a sum rule for the matrix elements of observables lydzba_swietek_24 , which we impose as
@@ II Scenarios for the ETH breakdown
where $\cal D$ is the Hilbert space dimension and $O_{nm}\equiv\langle n|\hat{O}|m\rangle$, see Appendix A for details.
@@
We note that recent studies based on the norms of adiabatic gauge potentials contributed valuable insights into the structure of the off-diagonal matrix elements in the vicinity of the ergodicity breakdown pandey_claeys_20 . In particular, they suggested that the softening of fluctuations of matrix elements at low $\omega$ is s smooth process in the vicinity of the breakdown leblond_sels_21 ; kim_polkovnikov_24 . However, a clear distinction between the scenarios (a) and (b) has to our knowledge not yet been established.
Here we put forth arguments in favor of scenario (b). Our main conjecture is that the fluctuations of the diagonal and low-$\omega$ off-diagonal matrix elements soften when approaching the ergodicity boundary, such that the fluctuating part of the ETH ansatz in Eq. (1) should acquire $\omega$-dependence, $\rho(\bar{E})^{-1/2}\to\Sigma(\bar{E},\omega,L)$. Hence, the difference of our scenario with respect to the conventional ETH is that the $L$-dependence of fluctuations, expressed via the density of states $\rho(\bar{E})$, and the $\omega$-dependence of the off-diagonal matrix elements, do not decouple as in Eq. (1). In the low-$\omega$ regime $\omega\to 0$, we express the softening of fluctuations as
@@
with $2<\eta<\infty$ in the fading ergodicity regime. We propose that $\eta$ is determined by the ratio of the Thouless and Heisenberg energy, and when the Thouless time is sufficiently large, the divergence of $\eta$ signals proximity of the ergodicity breaking phase transition. This scenario is sketched in Fig. 1.
@@
III Softening of ETH at small energies
We now present an argument supporting Eq. (3), i.e., a scenario that naturally interpolates between the conventional ETH ($\eta=2$) and its breakdown in the non-ergodic limit ($\eta\to\infty$). The argument is based on the $L$-dependence of the off-diagonal matrix elements at low $\omega$ for the targeted observable $\hat{O}$. For simplicity, we model the coarse-grained off-diagonal matrix elements at low $\omega$ with a Lorentzian function. This specific functional form is not crucial for the derivation of our main result, however, since we are approaching the ergodicity boundary from the ergodic side, it is important that the function is a constant in the $\omega\to 0$ limit. We consider
@@ III Softening of ETH at small energies
Importantly, the Thouless energy $\Gamma$ depends on $L$ in most physical systems (it vanishes as $L\to\infty$). Consequently, also the scaled off-diagonal matrix elements $\overline{|O_{nm}|^{2}}\rho$ are expected to depend on $L$. This is a known property, which is usually interpreted via the $L$-dependence of $f(\bar{E},\omega)$, see Eq. (1), at $\omega\ll\Gamma$ dalessio_kafri_16 ; luitz_barlev_16 ; brenes_leblond_20 ; richter_dymarsky_20 ; schoenle_jansen_21 ; wang_lamann_22 ; Sugimoto2022 ; balachandran_santos_23 . However, in many cases this $L$-dependence is polynomial, and it is hence subleading when compared to the exponentially increasing $\rho$.
@@
IV Example: Quantum sun model
We now present analytical and numerical evidence that supports the above scenario in a toy model that hosts an ergodicity breaking phase transition in the thermodynamic limit, i.e., the quantum sun model of the avalanche theory deroeck_huveneers_17 ; luitz_huveneers_17 ; suntajs_vidmar_22 . The Hamiltonian is defined as
$\hat{H}=\hat{H}_{\rm dot}+\sum_{j=1}^{L}\alpha^{u_{j}}\hat{S}_{n(j)}^{x}\hat{S}_{j}^{x}+\sum_{j=1}^{L}h_{j}\hat{S}_{j}^{z},$ (7)
where $\hat{H}_{\rm dot}$ is a $2^{N}\times 2^{N}$ random matrix drawn from the Gaussian orthogonal ensemble (GOE) describing all-to-all interacting particles within an ergodic quantum dot (we set $N=3$). The second term defines the coupling between a spin $j$ outside the dot ($j=1,...,L)$ and a randomly selected spin $n(j)$ within the dot, with $\alpha$ acting as the tuning parameter of the ergodicity breaking phase transition, and $u_{j}\propto j$. Details of the model implementation and the choice of model parameters are given in Appendix B.
@@ IV Example: Quantum sun model
Below we numerically test the above predictions. We focus on the observable $\hat{O}=\hat{S}_{L}^{z}$, i.e., the $z$-component of the most weakly coupled spin to the ergodic quantum dot. The physical motivation for selecting this observable is based on understanding that observables, which measure properties of spins that are most weakly coupled to the ergodic quantum dot, are most sensitive to ergodicity breaking suntajs_hopjan_24 , and hence their relaxation time can be associated with the Thouless time. Results for other observables are shown in Appendix C.
IV.1 Fluctuations of matrix elements
@@
In Fig. 2, we show the $\omega$-dependence of the coarse-grained off-diagonal matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}$, where $(S_{L}^{z})_{nm}\equiv\langle n|\hat{S}_{L}^{z}|m\rangle$, and the coarse graining (followed by the averaging over Hamiltonian realizations) is carried out over a narrow window above the target $\omega$. The scaled matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}\rho$ vs $\omega$ exhibit accumulation of spectral weight at low $\omega$ and its depletion at large $\omega$, see the arrows in the main panel of Fig. 2. Remarkably, the results in Fig. 2 are shown for $\alpha=0.86$, for which the nearest level-spacing statistics complies with the GOE predictions, see Fig. 7(a) and Appendix B. This suggests that the softening of the fluctuations of the matrix elements emerges in the fading ergodicity regime, in which the short-range statistics are still GOE-like. Another perspective on the softening of the low-$\omega$ matrix elements is shown in the inset of Fig. 2, in which we show the scaled matrix elements $\overline{|(S_{L}^{z})_{nm}|^{2}}\rho\Gamma$ vs $\omega/\Gamma$, exhibiting a reasonably good collapse of the results for all system sizes under investigation.
@@
Figure 3: Scaling of fluctuations of matrix elements. (a), (b) Eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, $\delta_{n}\equiv|(\hat{S}^{z}_{L})_{n+1}-(\hat{S}^{z}_{L})_{n}|$. We show the average and the maximal outliers in (a) and (b), respectively, see also Appendix A for the details on the intervals of averaging. (c) Low-$\omega$ off-diagonal matrix elements, taken from a narrow energy window around the target $\omega=\sqrt{\Gamma\Delta}$. Dashed lines in (a)-(c) are the two-parameter fits of the function $a_{0}2^{-L/\eta}$ with the values of $\eta$ shown next to them. (d) Fluctuation exponents $\eta$ from Eq. (3) as a function of $\alpha$. The solid line is a fit of $b_{0}\eta^{*}$ to the results for the off-diagonal matrix elements, where $b_{0}=1.12$ is the fitting parameter and $\eta^{*}$ is the function from Eq. (10), in which $\tilde{\alpha}_{c}$ is replaced by a fitting parameter $\alpha_{c}$. We get $\alpha_{c}=0.745$, see the vertical dashed-dotted line, which is very close to the values of transition point obtained from other indicators such as the average gap ratio shown in Fig. 7(b) of Appendix B.
@@
A quantitative analysis of the fluctuations of the matrix elements in carried out in Fig. 3. We study the average eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, see Fig. 3(a), and the fluctuations of the low-$\omega$ off-diagonal matrix elements measured at $\omega=\sqrt{\Delta\Gamma}$, see Fig. 3(c). They both exhibit a fast decay deep in the ergodic regime (at $\alpha$ close to $\alpha=1$), while the decay becomes much slower when the ergodicity breaking transition point is approached, $\alpha\to\alpha_{c}$. The extracted fluctuation exponents $\eta$ as functions of $\alpha$ are shown as symbols in Fig. 3(d). They behave very similarly for the diagonal and the low-$\omega$ off-diagonal matrix elements, and remarkably, are well described by the solid line that corresponds to the analytical prediction from Eq. (10). Moreover, Fig. 3(b) shows that even the maximal outliers of the diagonal matrix elements decay to zero with increasing $L$. This property suggest that the system is still ergodic, and makes a clear distinction to non-ergodic systems such as integrable interacting systems in which the maximal outliers do not decay with system size biroli_lauchli_10 ; leblond_mallayya_19 ; lydzba_mierzejewski_23 .
While the focus in the main text is on the fluctuations of matrix elements, we also study their distributions in Appendix D. There, we show evidence that the distributions appear to be close to a normal distribution in nearly entire ergodic phase.
IV.2 Quantum dynamics
We now demonstrate the emergence of ergodicity, and its breakdown, from the dynamical perspective. Our main goal is to show that the fading ergodicity regime can still be characterized as ergodic despite the breakdown of the conventional ETH. On the other hand, we show evidence that the breakdown of ergodicity occurs when the fluctuation exponent diverges, $\eta\to\infty$.
We study the dynamics of the expectation value of $\hat{S}_{L}^{z}$ after a sudden quantum quench. For a single Hamiltonian realization $\hat{H}^{(\mu)}$, it is defined as
$\mathcal{Q}(t)^{(\mu)}=\langle\psi_{0}|\hat{S}_{L}^{z}(t)|\psi_{0}\rangle\;,$ (11)
where $\hat{S}_{L}^{z}(t)=\exp(i\hat{H}^{(\mu)}t)\hat{S}_{L}^{z}\exp(-i\hat{H}^{(\mu)}t)$. The initial state $|\psi_{0}\rangle$ is a product state in the middle of the spectrum, see Appendix E for details. To study whether the system thermalizes or not, we first define the expectation values of $\hat{S}_{L}^{z}$ in the diagonal ensemble, $\mathcal{Q}_{\rm DE}^{(\mu)}$, and in the microcanonical ensemble, $\mathcal{Q}^{(\mu)}_{\rm ME}$, see Eqs. (26) and (27) for definitions. This allows us to compare the long-time value after a quantum quench to the microcanonical ensemble prediction via the quantity
$\Delta\mathcal{Q}_{\infty}={\rm Avr}_{\mu}\left\{|\mathcal{Q}_{\rm DE}^{(\mu)}-\mathcal{Q}^{(\mu)}_{\rm ME}|\right\}\;,$ (12)
where ${\rm Avr}_{\mu}\{...\}$ corresponds to the average over Hamiltonian realizations.